# Gravitational Waves Emission in Quadratic Gravity: longitudinal modes, angular momentum emission, and positivity of the radiated power

### v1

@@ II.1.1 The energy-momentum tensor of the GW in quadratic gravity
The energy-momentum tensor $t_{\mu\nu}$ in Eq. (32) can be compared to the $t_{\mu\nu}$ computed in Ref. [40]. This reference studies gravitational waves from inspiraling black holes in quadratic gravity; the modified gravity theory therein is the same as the one explored herein, thus one would expect consistency of the results. In fact, Eq. (32) differs from Eq. (53) of Ref. [40] by the term $\left\langle\partial_{\mu}\tilde{h}^{\alpha\beta}\partial_{\nu}\Psi_{\alpha\beta}\right\rangle$, which is present in the latter. However, it was shown in Ref. [40] that this additional term vanishes. Therefore, the geometrical approach to computing $t_{\mu\nu}$ adopted by Ref. [40] is in complete accordance with the field-theory approach adopted in this work.
@@ II.1.2 The angular momentum of the GW in quadratic gravity
$A_{kl}^{i}=\delta^{ik}x^{l}-\delta^{il}x^{k}\qquad\text{and}\qquad A_{kl}^{0}=0.$ (35)
A quick check of Eqs. (17) and (28)–(30) shows that the rotation affects not only the coordinates but also the fields $\tilde{h}_{\alpha\beta}$, $\Psi_{\alpha\beta}$ and $\Phi$. The field $\Phi$ is invariant under rotations. The fields $\tilde{h}_{\alpha\beta}$ and $\Psi_{\alpha\beta}$ may be decomposed into scalar, vector, tensor irreducible quantities.
@@
$B_{0i}^{\prime}\left(x^{\prime}\right)$ $=$ $\mathcal{R}_{i}^{\hphantom{i}j}B_{0j}\Rightarrow$ (37)
$B_{0i}^{\prime}\left(x^{\prime}\right)$ $=$ $B_{0i}+\sum_{k<l}\omega^{kl}\left(\delta_{ik}B_{0l}-\delta_{il}B_{0k}\right);$
$\bar{B}_{ij}^{\prime}$ $=$ $\mathcal{R}_{i}^{\hphantom{i}k}\mathcal{R}_{j}^{\hphantom{i}l}\bar{B}_{kl}\Rightarrow$ (38)
$\bar{B}_{ij}^{\prime}$ $=$ $\bar{B}_{ij}+\sum_{k<l}\omega^{kl}\left(\delta_{kj}\bar{B}_{il}-\delta_{jl}\bar{B}_{ik}+\delta_{ik}\bar{B}_{lj}-\delta_{il}\bar{B}_{kj}\right).$
Eqs. (36)–(38) allows us to identify the quantities $F_{\alpha\beta,a}$ in (17) describing field rotations. In fact, we have $a=\left[ik\right]$ with $\left(i,j\right)=1,2,3$,
@@
By expliciting the average instructions, it is possible to demonstrate that $\left\langle\left(CT\right)_{l}\right\rangle$ is zero. In fact, integrations by parts in the average of the first term in Eq. (53) lead to surface terms—which can be neglected —and to the expression:
@@
$\left.-\alpha\partial_{t}\square\bar{h}^{ij}-\alpha\partial_{t}\square\bar{\Psi}^{ij}\right]F_{ij,kl}^{\tilde{h}},$
@@
$\left.-2\alpha\partial_{t}\square\tilde{h}^{0i}-2\alpha\partial_{t}\square\Psi^{0i}\right]F_{0i,kl}^{\Psi}$
$+\frac{1}{2\kappa c}\left[\frac{1}{2}\partial_{t}\bar{\Psi}^{ij}+\frac{1}{2}\partial_{t}\bar{h}^{ij}\right.$ (56)
$-\left.\alpha\partial_{t}\square\bar{h}^{ij}-\alpha\partial_{t}\square\bar{\Psi}^{ij}\right]F_{ij,kl}^{\Psi}.$
@@
$\left(s_{kl}^{0}\right)^{\Psi}$ $=-\frac{1}{\kappa c}\left[\partial_{t}\Psi^{0i}+2c\partial^{i}\Phi\right]F_{0i,kl}^{\Psi}$ (58)
$-\frac{1}{4\kappa c}\left[\partial_{t}\Psi^{ij}-\partial_{t}\tilde{h}^{ij}\right]F_{ij,kl}^{\Psi}.$ (59)
@@
$\left.+2\Psi_{0[l}\dot{\Psi}_{k]0}+\left(CT\right)_{s_{1}}-4c\left(CT\right)_{s_{2}}\right]$
@@
Subtracting (64) and (65) and taking the average, yields:
@@
After these necessary intermediate calculations, we substitute Eqs. (43), (52), and (II.1.2) into Eq. (42) to obtain the final form of the total angular momentum for GW in quadratic gravity:
$J^{i}=$ $\frac{c^{2}}{32\pi G}\int d^{3}x\left[\epsilon^{ikl}\left(-\left\langle\dot{h}_{ab}^{TT}x_{k}\partial_{l}\tilde{h}_{ab}^{TT}\right\rangle+2\left\langle h_{ak}^{TT}\dot{h}_{al}^{TT}\right\rangle\right.\right.$
$\qquad\qquad+\left\langle\dot{\Psi}^{\alpha\beta}x_{k}\partial_{l}\Psi_{\alpha\beta}\right\rangle-2\left\langle\Psi_{ak}\dot{\Psi}_{al}\right\rangle$
$\left.\left.+4\left\langle\Psi_{0k}\dot{\Psi}_{0l}\right\rangle-6\left\langle\dot{\Phi}x_{k}\partial_{l}\Phi\right\rangle\right)\right].$
The first two terms of the integrand reproduce the result known from general relativity [47]. The last four are extra terms related to the spin-2 and spin-0 massive fields.
@@ II.1.3 Energy and angular momentum radiated
We are interested in computing the energy and angular momentum carried away from the source by the gravitational waves. The gravitational radiation at a sufficiently distant point from the source (radiation zone) is structurally given by [40, 48]:
@@
$\left.-\left\langle\partial_{0}\Psi_{\left(n\right)}^{\alpha\beta}\partial_{r}\Psi_{\alpha\beta}^{\left(m\right)}\right\rangle+6\left\langle\partial_{0}\Phi_{\left(n\right)}\partial_{r}\Phi^{\left(m\right)}\right\rangle\right].$
@@
$\sum_{m,n}\left\langle\partial_{0}f_{\left(n\right)}^{i}\partial_{r}f_{i}^{\left(m\right)}\right\rangle$ $=\sum_{m,n}\left\langle\partial_{0}f_{\left(n\right)}^{i}\partial_{r}f_{i}^{\left(m\right)}\right\rangle\delta_{\left(m\right)}^{\left(n\right)}$ (74)
@@
$\frac{dE}{dt}$ $=$ $\frac{c^{3}r^{2}}{32\pi G}\sum_{n}\int d\Omega\left[\left\langle\dot{h}_{\left(n\right)}^{\alpha\beta}\dot{h}_{\alpha\beta}^{\left(n\right)}\right\rangle\right.$ (78)
$\left.-q_{n}^{\Psi}\left\langle\dot{\Psi}_{\left(n\right)}^{\alpha\beta}\dot{\Psi}_{\alpha\beta}^{\left(n\right)}\right\rangle\Theta_{n}^{\Psi}+6q_{n}^{\Phi}\left\langle\dot{\Phi}_{\left(n\right)}^{2}\right\rangle\Theta_{n}^{\Phi}\right],$
@@
$q_{n}^{\Psi}\equiv\sqrt{1-\left(\frac{m_{\Psi}c}{\omega_{n}}\right)^{2}},$ (79)
$q_{n}^{\Phi}\equiv\sqrt{1-\left(\frac{m_{\Phi}c}{\omega_{n}}\right)^{2}},$ (80)
@@
$\Theta_{n}^{\Psi}\equiv\Theta\left(\omega_{n}-m_{\Psi}c\right),$
$\Theta_{n}^{\Phi}\equiv\Theta\left(\omega_{n}-m_{\Phi}c\right),$ (81)
@@
The equation for the radiated angular momentum can be obtained from the conservation of the current $j_{kl}^{\mu}$ given by Eq. (43). However, it is simpler to perform a heuristic construction based on the total angular momentum [47]. In fact, the following expression for the angular momentum density $j^{i}/c$ is inferred from the structure of Eq. (II.1.2):
@@
$\frac{j_{\Psi}^{i\left(n\right)}}{c}$ $=\frac{r^{2}c^{2}}{32\pi G}\epsilon^{ikl}\left[\left\langle\dot{\Psi}_{\left(n\right)}^{\alpha\beta}x_{k}\partial_{l}\Psi_{\alpha\beta}^{\left(n\right)}\right\rangle-2\left\langle\Psi_{ak}^{\left(n\right)}\dot{\Psi}_{al}^{\left(n\right)}\right\rangle\right.$ (84)
$\left.+4\left\langle\Psi_{0k}^{\left(n\right)}\dot{\Psi}_{0l}^{\left(n\right)}\right\rangle\right],$
@@
The radiated gravitational wave of frequency $\omega_{n}$ is built from the various modes $h_{\alpha\beta}^{\left(n\right)}$, $\Psi_{\alpha\beta}^{\left(n\right)}$, and $\Phi^{\left(n\right)}$ that propagate with group velocities $v_{g}^{\left(n\right)}$: [^6] Section VI of Ref. [40] provides details on the group velocity of massive modes.
@@ III Quadrupole approximation
The solutions to the above equations in a region devoid of sources and in the quadrupole approximation are [47, 40]:
@@
$\Psi_{ij}\left(\mathbf{x},t\right)$ $=-\frac{\kappa}{4\pi r}\ddot{M}_{ij}\left(t_{r}\right)$ (92)
@@
Here $\Lambda_{ij,kl}$ is the projection tensor in the TT gauge [47], $t_{r}=t-\frac{r}{c}$ is the retarded time, $\zeta\equiv t_{r}-\bar{t}_{r}$, and $n_{i}$ points in the direction of propagation of the wave. The quadrupole mass moments are computed from the expressions
@@
with the Bessel function of first type $J_{1}$ and, again, $X=\Psi$,$\Phi$. We refer the interested reader to Ref. [40] for further details.
@@ III.1 GWs propagation in a generic direction
The massive spin-2 mode has five degrees of freedom. Aiming for a description similar to that in GR, we define [40]:
$\Psi_{+}=\frac{\Psi_{11}-\Psi_{22}}{2},\qquad\Psi_{\times}=\Psi_{12}\text{,}$
$\Psi_{B}=\Psi_{13}\text{, }\Psi_{C}=\Psi_{23}\qquad\text{and}\quad\Psi_{D}=\Psi_{33},$ (101)
@@
where $\theta$ is the polar angle and $\phi$ is the azimuth angle [47]. The rotation matrix connecting the directions $n_{i}$ and $n_{i}^{\prime}$ is:
@@
$M_{ij}$ $=R_{ia}R_{jb}M_{ab}^{\prime}\Rightarrow$ (114)
$\mathbf{M}$ $=\mathbf{RM}^{\prime}\mathbf{R}^{T}\Rightarrow$ (115)
@@
    $-\ddot{M}_{11}\cos^{2}\theta\sin^{2}\phi+\ddot{M}_{22}\sin^{2}\phi$
    $-\ddot{M}_{22}\cos^{2}\theta\cos^{2}\phi-\ddot{M}_{33}\sin^{2}\theta$
    $-\ddot{M}_{12}\left(1+\cos^{2}\theta\right)\sin\left(2\phi\right)$
@@
    $-2\ddot{M}_{13}\sin\theta\cos\phi$
    $\left.+2\ddot{M}_{23}\sin\theta\sin\phi\right],$ (120)
    $\Psi_{B}\left(t,\theta,\phi\right)$ $=-\frac{\kappa}{8\pi r}\left[1-\hat{O}_{\Psi}\right]\left[\left(\ddot{M}_{11}-\ddot{M}_{22}\right)\sin\theta\sin\left(2\phi\right)\right.$
    $+2\ddot{M}_{12}\sin\theta\cos\left(2\phi\right)+2\ddot{M}_{13}\cos\theta\cos\phi$ (121)
    $\left.-2\ddot{M}_{23}\cos\theta\sin\phi\right],$
@@
    $+\ddot{M}_{22}\cos^{2}\phi\sin\left(2\theta\right)$
    $-\ddot{M}_{33}\sin\left(2\theta\right)+\ddot{M}_{12}\sin\left(2\theta\right)\sin\left(2\phi\right)$
    $\left.+2\ddot{M}_{13}\cos\left(2\theta\right)\sin\phi+2\ddot{M}_{23}\cos\left(2\theta\right)\cos\phi\right],$ (122)
    and
    $\Psi_{D}$ $=-\frac{\kappa}{4\pi r}\left[1-\hat{O}_{\Psi}\right]\left[\ddot{M}_{11}\sin^{2}\phi\sin^{2}\theta\right.$
@@
    $+\ddot{M}_{12}\sin^{2}\theta\sin\left(2\phi\right)$
    $\left.+\ddot{M}_{13}\sin\left(2\theta\right)\sin\phi+\ddot{M}_{23}\sin\left(2\theta\right)\cos\phi\right].$ (124)
@@
$\bar{h}_{ij}^{\left(n\right)}$ $=\left(\begin{array}[]{ccc}h_{+}^{\left(n\right)}+\Psi_{+}^{\left(n\right)}&h_{\times}^{\left(n\right)}+\Psi_{\times}^{\left(n\right)}&0\\ h_{\times}^{\left(n\right)}+\Psi_{\times}^{\left(n\right)}&-\left(h_{+}^{\left(n\right)}+\Psi_{+}^{\left(n\right)}\right)&0\\ 0&0&0\end{array}\right)+\left(\begin{array}[]{ccc}\Phi^{\left(n\right)}&0&0\\ 0&\Phi^{\left(n\right)}&0\\ 0&0&\Phi^{\left(n\right)}\end{array}\right)+\frac{1}{2}\left(\begin{array}[]{ccc}\left(\left(q_{n}^{\Psi}\right)^{2}-1\right)\Psi_{D}^{\left(n\right)}&0&\Psi_{B}^{\left(n\right)}\\ 0&\left(\left(q_{n}^{\Psi}\right)^{2}-1\right)\Psi_{D}^{\left(n\right)}&\Psi_{C}^{\left(n\right)}\\ \Psi_{B}^{\left(n\right)}&\Psi_{C}^{\left(n\right)}&2\Psi_{D}^{\left(n\right)}\end{array}\right).$ (134)

### v2

@@ II.1.1 The energy-momentum tensor of the GW in quadratic gravity
The energy-momentum tensor $t_{\mu\nu}$ in Eq. (32) can be compared to the $t_{\mu\nu}$ computed in Ref. Alves et al. (2023). This reference studies gravitational waves from inspiraling black holes in quadratic gravity; the modified gravity theory therein is the same as the one explored herein, thus one would expect consistency of the results. In fact, Eq. (32) differs from Eq. (53) of Ref. Alves et al. (2023) by the term $\left\langle\partial_{\mu}\tilde{h}^{\alpha\beta}\partial_{\nu}\Psi_{\alpha\beta}\right\rangle$, which is present in the latter. However, it was shown in Ref. Alves et al. (2023) that this additional term vanishes. Therefore, the geometrical approach to computing $t_{\mu\nu}$ adopted by Ref. Alves et al. (2023) is in complete accordance with the field-theory approach adopted in this work. Here, there is a cancelation of the second term on the right-hand side (r.h.s.) of Eq. (28) with the first term on the r.h.s. of Eq. (29) on average after we neglect topological terms. The physical difference between the field-theoretical approach and the geometrical approach is precisely due to these topological terms. Because these terms are discarded in the bulk, the results of both approaches converge.
@@ II.1.2 The angular momentum of the GW in quadratic gravity
$A_{kl}^{i}=\delta^{i}_{k}x_{l}-\delta^{i}_{l}x_{k}\qquad\text{and}\qquad A_{kl}^{0}=0.$ (35)
A quick check of Eqs. (17) and (28)–(30) shows that the rotation affects not only the coordinates but also the fields $\tilde{h}_{\alpha\beta}$, $\Psi_{\alpha\beta}$ and $\Phi$. The field $\Phi$ is invariant under rotations. The fields $\tilde{h}_{\alpha\beta}$ and $\Psi_{\alpha\beta}$ may be decomposed into scalar, vector, and tensor irreducible quantities.
@@
$B_{0i}^{\prime}\left(x^{\prime}\right)$ $=\mathcal{R}_{i}^{\hphantom{i}j}B_{0j}$
$=B_{0i}+\sum_{k<l}\omega^{kl}\left(\delta_{ik}B_{0l}-\delta_{il}B_{0k}\right);$ (37)
$\bar{B}_{ij}^{\prime}$ $=\mathcal{R}_{i}^{\hphantom{i}k}\mathcal{R}_{j}^{\hphantom{i}l}\bar{B}_{kl}$
$=\bar{B}_{ij}+\sum_{k<l}\omega^{kl}\left(\delta_{kj}\bar{B}_{il}-\delta_{jl}\bar{B}_{ik}+\delta_{ik}\bar{B}_{lj}-\delta_{il}\bar{B}_{kj}\right).$ (38)
Eqs. (36)–(38) allow us to identify the quantities $F_{\alpha\beta,a}$ in (17) describing field rotations. In fact, we have $a=\left[ik\right]$ with $\left(i,j\right)=1,2,3$, as follows:
@@
By expliciting the average instructions, it is possible to demonstrate that $\left\langle\left(CT\right)_{l}\right\rangle$ is zero. In fact, integrations by parts on average of the first term in Eq. (53) lead to surface terms—which can be neglected —and to the expression:
@@
$\left.-\alpha\partial_{t}\square\bar{h}^{ij}-\alpha\partial_{t}\square\bar{\Psi}^{ij}\right]F_{ij,kl}^{\tilde{h}},$ (55)
@@
$\qquad\left.-2\alpha\partial_{t}\square\tilde{h}^{0i}-2\alpha\partial_{t}\square\Psi^{0i}\right]F_{0i,kl}^{\Psi}$
$+\frac{1}{2\kappa c}\left[\frac{1}{2}\partial_{t}\bar{\Psi}^{ij}+\frac{1}{2}\partial_{t}\bar{h}^{ij}\right.$
$\qquad-\left.\alpha\partial_{t}\square\bar{h}^{ij}-\alpha\partial_{t}\square\bar{\Psi}^{ij}\right]F_{ij,kl}^{\Psi},$ (56)
@@
$\left(s_{kl}^{0}\right)^{\Psi}$ $=-\frac{1}{\kappa c}\left[\partial_{t}\Psi^{0i}+2c\partial^{i}\Phi\right]F_{0i,kl}^{\Psi}$
$\quad\,-\frac{1}{4\kappa c}\left[\partial_{t}\Psi^{ij}-\partial_{t}\tilde{h}^{ij}\right]F_{ij,kl}^{\Psi}.$ (58)
@@
$\left.+2\Psi_{0[l}\dot{\Psi}_{k]0}+\left(CT\right)_{s_{1}}-4c\left(CT\right)_{s_{2}}\right]$ (59)
@@
Subtracting (63) and (64) and taking the average yields
@@
After these necessary intermediate calculations, we substitute Eqs. (43), (52), and (59) into Eq. (42) to obtain the final form of the total angular momentum for GW in quadratic gravity:
$J^{i}$ $=\frac{c^{2}}{32\pi G}\int d^{3}x\left[\epsilon^{ikl}\left(-\left\langle\dot{h}_{ab}^{TT}x_{k}\partial_{l}\tilde{h}_{ab}^{TT}\right\rangle\right.\right.$
$+2\left\langle h_{ak}^{TT}\dot{h}_{al}^{TT}\right\rangle+\left\langle\dot{\Psi}^{\alpha\beta}x_{k}\partial_{l}\Psi_{\alpha\beta}\right\rangle-2\left\langle\Psi_{ak}\dot{\Psi}_{al}\right\rangle$
$\left.\left.+4\left\langle\Psi_{0k}\dot{\Psi}_{0l}\right\rangle-6\left\langle\dot{\Phi}x_{k}\partial_{l}\Phi\right\rangle\right)\right].$ (66)
The first two terms of the integrand reproduce the result known from general relativity Maggiore (2007). The last four are extra terms related to the spin-2 and spin-0 massive fields.
@@ II.1.3 Energy and angular momentum radiated
We are interested in computing the energy and angular momentum carried away from the source by the gravitational waves. The gravitational radiation at a sufficiently distant point from the source (radiation zone) is structurally given by Refs. Alves et al. (2023); Vilhena et al. (2021):
@@
$\left.-\left\langle\partial_{0}\Psi_{\left(n\right)}^{\alpha\beta}\partial_{r}\Psi_{\alpha\beta}^{\left(m\right)}\right\rangle+6\left\langle\partial_{0}\Phi_{\left(n\right)}\partial_{r}\Phi^{\left(m\right)}\right\rangle\right].$ (71)
@@
$\sum_{m,n}\left\langle\partial_{0}f_{\left(n\right)}^{i}\partial_{r}f_{i}^{\left(m\right)}\right\rangle$ $=\sum_{m,n}\left\langle\partial_{0}f_{\left(n\right)}^{i}\partial_{r}f_{i}^{\left(m\right)}\right\rangle\delta_{\left(m\right)}^{\left(n\right)}$
@@
$\frac{dE}{dt}$ $=\frac{c^{3}r^{2}}{32\pi G}\sum_{n}\int d\Omega\left[\left\langle\dot{h}_{\left(n\right)}^{\alpha\beta}\dot{h}_{\alpha\beta}^{\left(n\right)}\right\rangle\right.$
$\left.-q_{n}^{\Psi}\left\langle\dot{\Psi}_{\left(n\right)}^{\alpha\beta}\dot{\Psi}_{\alpha\beta}^{\left(n\right)}\right\rangle\Theta_{n}^{\Psi}+6q_{n}^{\Phi}\left\langle\dot{\Phi}_{\left(n\right)}^{2}\right\rangle\Theta_{n}^{\Phi}\right],$ (75)
@@
$q_{n}^{\Psi}\equiv\sqrt{1-\left(\frac{m_{\Psi}c}{\omega_{n}}\right)^{2}},\quad\rm{and}\quad q_{n}^{\Phi}\equiv\sqrt{1-\left(\frac{m_{\Phi}c}{\omega_{n}}\right)^{2}},$ (76)
@@
$\Theta_{n}^{\Psi}\equiv\Theta\left(\omega_{n}-m_{\Psi}c\right),\quad\rm{and}\quad\Theta_{n}^{\Phi}\equiv\Theta\left(\omega_{n}-m_{\Phi}c\right),$ (77)
@@
The equation for the radiated angular momentum can be obtained from the conservation of the current $j_{kl}^{\mu}$ given by Eq. (43). However, it is simpler to perform a heuristic construction based on the total angular momentum Maggiore (2007). In fact, the following expression for the angular momentum density $j^{i}/c$ is inferred from the structure of Eq. (66):
@@
$\frac{j_{\Psi}^{i\left(n\right)}}{c}$ $=\frac{r^{2}c^{2}}{32\pi G}\epsilon^{ikl}\left[\left\langle\dot{\Psi}_{\left(n\right)}^{\alpha\beta}x_{k}\partial_{l}\Psi_{\alpha\beta}^{\left(n\right)}\right\rangle\right.$
$\left.-2\left\langle\Psi_{ak}^{\left(n\right)}\dot{\Psi}_{al}^{\left(n\right)}\right\rangle+4\left\langle\Psi_{0k}^{\left(n\right)}\dot{\Psi}_{0l}^{\left(n\right)}\right\rangle\right],$ (80)
@@
The radiated gravitational wave of frequency $\omega_{n}$ is built from the various modes $h_{\alpha\beta}^{\left(n\right)}$, $\Psi_{\alpha\beta}^{\left(n\right)}$, and $\Phi^{\left(n\right)}$ that propagate with group velocities $v_{g}^{\left(n\right)}$: [^6] Section VI of Ref. Alves et al. (2023) provides details on the group velocity of massive modes.
@@ III Quadrupole approximation
The solutions to the above equations in a region devoid of sources and in the quadrupole approximation are Maggiore (2007); Alves et al. (2023):
@@
$\Psi_{ij}\left(\mathbf{x},t\right)$ $=-\frac{\kappa}{4\pi r}\ddot{M}_{ij}\left(t_{r}\right)$
@@
Here $\Lambda_{ij,kl}$ is the projection tensor in the TT gauge Maggiore (2007), $t_{r}=t-\frac{r}{c}$ is the retarded time, $\zeta\equiv t_{r}-\bar{t}_{r}$, and $n_{i}$ points in the direction of propagation of the wave. The quadrupole mass moments are computed from the expressions
@@
with the Bessel function of first type $J_{1}$ and, again, $X=\Psi$,$\Phi$. We refer the interested reader to Ref. Alves et al. (2023) for further details.
@@ III.1 GWs propagation in a generic direction
The massive spin-2 mode has five degrees of freedom. Aiming for a description similar to that in GR, we define Alves et al. (2023):
$\Psi_{+}=\frac{\Psi_{11}-\Psi_{22}}{2},\quad\Psi_{\times}=\Psi_{12},$
$\Psi_{B}=\Psi_{13},\quad\Psi_{C}=\Psi_{23}\,\quad\text{and}\quad\Psi_{D}=\Psi_{33},$ (97)
@@
where $\theta$ is the polar angle and $\phi$ is the azimuth angle Maggiore (2007). The rotation matrix connecting the directions $n_{i}$ and $n_{i}^{\prime}$ is:
@@
$M_{ij}=R_{ia}R_{jb}M_{ab}^{\prime}\Rightarrow\mathbf{M}=\mathbf{RM}^{\prime}\mathbf{R}^{T}$
i.e.,
@@
    $-\ddot{M}_{11}\cos^{2}\theta\sin^{2}\phi+\ddot{M}_{22}\sin^{2}\phi-\ddot{M}_{22}\cos^{2}\theta\cos^{2}\phi$
    $-\ddot{M}_{33}\sin^{2}\theta-\ddot{M}_{12}\left(1+\cos^{2}\theta\right)\sin\left(2\phi\right)$
@@
    $\left.-2\ddot{M}_{13}\sin\theta\cos\phi+2\ddot{M}_{23}\sin\theta\sin\phi\right],$ (114)
    $\Psi_{B}\left(t,\theta,\phi\right)$ $=-\frac{\kappa}{8\pi r}\left[1-\hat{O}_{\Psi}\right]\left[\ddot{M}_{11}\sin\theta\sin\left(2\phi\right)\right.$
    $-\ddot{M}_{22}\sin\theta\sin\left(2\phi\right)+2\ddot{M}_{12}\sin\theta\cos\left(2\phi\right)$
    $\left.+2\ddot{M}_{13}\cos\theta\cos\phi-2\ddot{M}_{23}\cos\theta\sin\phi\right],$ (115)
@@
    $+\ddot{M}_{22}\cos^{2}\phi\sin\left(2\theta\right)-\ddot{M}_{33}\sin\left(2\theta\right)$
    $+\ddot{M}_{12}\sin\left(2\theta\right)\sin\left(2\phi\right)+2\ddot{M}_{13}\cos\left(2\theta\right)\sin\phi$
    $\left.+2\ddot{M}_{23}\cos\left(2\theta\right)\cos\phi\right],$ (116)
    $\Psi_{D}\left(t,\theta,\phi\right)$ $=-\frac{\kappa}{4\pi r}\left[1-\hat{O}_{\Psi}\right]\left[\ddot{M}_{11}\sin^{2}\phi\sin^{2}\theta\right.$
@@
    $\left.+\ddot{M}_{23}\sin\left(2\theta\right)\cos\phi\right].$ (118)
@@
$\bar{h}_{ij}^{\left(n\right)}=\left(\begin{array}[]{ccc}h_{+}^{\left(n\right)}+\Psi_{+}^{\left(n\right)}&h_{\times}^{\left(n\right)}+\Psi_{\times}^{\left(n\right)}&0\\ h_{\times}^{\left(n\right)}+\Psi_{\times}^{\left(n\right)}&-\left(h_{+}^{\left(n\right)}+\Psi_{+}^{\left(n\right)}\right)&0\\ 0&0&0\end{array}\right)$ (122)
$+\left(\begin{array}[]{ccc}\Phi^{\left(n\right)}&0&0\\ 0&\Phi^{\left(n\right)}&0\\ 0&0&\Phi^{\left(n\right)}\end{array}\right)$ (126)
$+\frac{1}{2}\left(\begin{array}[]{ccc}\left(\left(q_{n}^{\Psi}\right)^{2}-1\right)\Psi_{D}^{\left(n\right)}&0&\Psi_{B}^{\left(n\right)}\\ 0&\left(\left(q_{n}^{\Psi}\right)^{2}-1\right)\Psi_{D}^{\left(n\right)}&\Psi_{C}^{\left(n\right)}\\ \Psi_{B}^{\left(n\right)}&\Psi_{C}^{\left(n\right)}&2\Psi_{D}^{\left(n\right)}\end{array}\right).$ (130)